OpenStudy (anonymous):

solve the compound inequality 5>-2x+4 or 9<=-2x+4

OpenStudy (anonymous):

Start by solving each inequality separately.

OpenStudy (asadkarim7):

5 - 4 > -2x x < -1/2 9 -4 <= -2x - 5 /2 <= x

myininaya (myininaya):

5-4>-2x 1>-2x of we divide(or multiply by a negative on both sides) then flip the inequality -1/2 <x now you try the second inequality let me know what you get

OpenStudy (anonymous):

(-1/2)

myininaya (myininaya):

lets see if thats right 9<=-2x+4 9-4<=-2x 5<=-2x -5/2>=x nope you get x<=-5/2 so we have x>-1/2 for the other inequality ~~~~* ~~~~ _____|______|______ -5/2 -1/2 they don't intersect so we write x is an element of (-inf,-5/2]U(-1/2,inf)

myininaya (myininaya):

oh asad i didn't realize you did the second one lol